if the sum of two consecutive integers is divided by 3, the quotient is 9. find the numbers
Found 2 solutions by Alan3354, mananth:
Answer by Alan3354(69443)

then the sum is 27.
27/2 = 13.5, the center.
--> 13 & 14

Answer by mananth(16949)

let one number be x
the other number will be x+1
.
x+x+1 / 3= 9
2x+1=9*3
2x=27-1
2x=26
x=13
the numbers are 13 and 14